# Forgetting Transformer: Softmax Attention with a Forget Gate

### v1

@@ Abstract
An essential component of modern recurrent sequence models is the forget gate. While Transformers do not have an explicit recurrent form, we show that a forget gate can be naturally incorporated into Transformers by down-weighting the unnormalized attention scores in a data-dependent way. We name this attention mechanism the Forgetting Attention and the resulting model the Forgetting Transformer (FoX). We show that FoX outperforms the Transformer on long-context language modeling, length extrapolation, and short-context downstream tasks, while performing on par with the Transformer on long-context downstream tasks. Moreover, it is compatible with the FlashAttention algorithm and does not require any positional embeddings. Several analyses, including the needle-in-the-haystack test, show that FoX also retains the Transformer’s superior long-context capabilities over recurrent sequence models such as Mamba-2, HGRN2, and DeltaNet. We also introduce a “Pro” block design that incorporates some common architectural components in recurrent sequence models and find it significantly improves the performance of both FoX and the Transformer. Our code is available at https://github.com/zhixuan-lin/forgetting-transformer.
@@ 1 Introduction
To address this question, we leverage an important fact: many recurrent sequence models with a forget gate can be written in a parallel linear attention form (Katharopoulos et al., 2020) analogous to softmax attention (Yang et al., 2023; Dao & Gu, 2024). In this parallel form, the forget gate mechanism translates into down-weighing the unnormalized attention scores in a data-dependent way. Our key insight is that this exact mechanism is also applicable to softmax attention. We name this attention mechanism the Forgetting Attention and the resulting model the Forgetting Transformer (FoX).
@@ 3 Forgetting Transformer
Our proposed model, the Forgetting Transformer (FoX), features a modified softmax attention mechanism with a forget gate. We name this attention mechanism the Forgetting Attention. Similar to the gated variant of linear attention introduced in the previous section, we first compute a scalar forget gate $f_{t}=\sigma({\bm{w}}_{f}^{\top}{\bm{x}}_{t}+b_{f})\in{\mathbb{R}}$ for each timestep $t$. The output of the attention is then
@@
where $F_{ij}=\prod_{l=j+1}^{i}f_{l}$ and $d_{ij}=\sum_{l=j+1}^{i}\log f_{l}$. This can be written in matrix form:
@@
Besides its natural connection to gated linear attention, the Forgetting Attention can also be seen as a data-dependent and learnable version of ALiBi (Press et al., 2021). ALiBi applies a data-independent bias $b_{ij}=-(i-j)m_{h}$ to the attention logits, where $m_{h}$ is a fixed slope specific to each head $h$. It is easy to show that ALiBi is equivalent to using a fixed, head-specific, and data-independent forget gate $f_{t}^{(h)}=\exp(-m_{h})$. In Section 4.5, we verify the superiority of data-dependent forget gates over ALiBi.
@@ Results
Figure 3: Visualization of the forget gate weight matrix ${\bm{F}}$ and the attention score matrix ${\bm{A}}$ from two heads in different layers. Since ${\bm{A}}$ is very sparse, we only show entries with scores larger than $0.5$. These results use FoX (Pro). More examples can be found in Appendix F.10.
@@ 6 Conclusion
We discuss several limitations of our work and potential future work. First, due to our limited computing resources, our main experiments only use models up to 760M parameters, 48B tokens, and a training context length of $16384$ tokens. Thus, an important direction for future work is to test FoX at larger scales. Second, we only consider causal sequence modeling in this work. It would be interesting to extend the Forgetting Attention to the non-causal case. Finally, we could potentially prune computation (e.g., KV-cache eviction) adaptively based on the forget gate values, which may greatly reduce training and inference costs.
@@ Appendix A Detailed FoX (Pro) layer computation
where $d_{ij}=\sum_{l=j+1}^{i}\log f_{l}$. We then apply the output normalization, output gate, and the final output projection:
@@
For the multi-head case, each head $h$ maintains an independent copy of the parameters and computes its output sequence $({\bm{y}}_{i}^{(h)})_{i=1}^{L}$ independently. ${\bm{y}}_{i}^{(h)}$’s from different heads are then summed together to obtain the final output ${\bm{y}}_{i}$. Note that similar to the standard Transformer, even though the computation and parameters of different heads are conceptually independent, the computation can be implemented equivalently by properly splitting/concatenating the intermediate vectors/weight matrices of different heads.
@@ B.2 Model parameters, estimated FLOPs, and throughput
We report the number of (non-embedding) parameters, estimated FLOPs, and training throughput in Table 6. For the recurrent models, we estimate FLOPs using their recurrent form for simplicity. Note that the FlashAttention kernels for FoX (Pro), Transformer (Pro), and FoX (LLaMA) are implemented in Triton by us on top of Flag Attention (FlagOpen, 2023) without significant optimization, while Transformer (LLaMA) uses the official FlashAttention implementation (Dao, 2023)) implemented in CUDA. Also, we did not focus on optimizing the efficiency of components such as QK-norm, KV-shift, and so on. We expect these four models to have similar throughput if they are properly optimized.
Since we use a long training context length compared to model size, recurrent models have a significant advantage in theoretical FLOPs due to their linear complexity. Though an exact FLOPs-matched comparison would be interesting, it will require recalibrating the scaling law for the long-context setting and is beyond the scope of this work.
@@ Appendix D Data-independent forget gate initialization
To understand our initialization for the fixed forget gate and the data-independent forget gate, we first define a function $T(b)=\frac{1}{-\log\sigma(b)}$. This function is defined such that $\sigma(b)^{T(b)}=1/e$ is always true (i.e., $T(b)$ is the timesteps needed to achieve a decay of $1/e$). We then initialize $b^{(h)}=b^{(h)}_{\text{init}}$ such that $T(b^{(h)}_{(\text{init})})=\exp(\log T_{\min}+(\log T_{\max}-\log T_{\min})\frac{h-1}{H-1})$, where $T_{\min}$ and $T_{\max}$ are hyperparameters and $H$ is the number of heads. For example, if $h=4$ and $(T_{\min},T_{\max})=(2,128)$, then we have $(T(b^{(1)}_{\text{init}}),T(b^{(2)}_{\text{init}}),T(b^{(3)}_{\text{init}}),T(b^{(4)}_{\text{init}}))=(2,8,32,128)$. As mentioned in the main text, A fixed forget gate with $(T_{\min},T_{\max})$ is equivalent to ALiBi with a minimum slope $\frac{1}{T_{\max}}$ and a maximum slope $\frac{1}{T_{\min}}$. For example, ALiBi with a maximum slope $\frac{1}{2}$ and a minimum slope $\frac{1}{256}$ (the default values in Press et al. (2021)) is equivalent to using a fixed forget gate with $(T_{\min},T_{\max})=(2,256)$.
We also tested this initialization for the data-dependent forget gate but did not find it useful, so we simply initialize $\{b^{(h)}\}_{h=1}^{H}$ to zero for the data-dependent forget gate. For the data-independent and fixed forget gate, zero initialization performs extremely poorly.
@@ F.10 Additional visualization of forget gate and attention score matrices
Figure 28: Visualization of the attention score matrix ${\bm{A}}$ from 16 heads in 4 different layers. These results use FoX (Pro).

### v2

@@ Abstract
An essential component of modern recurrent sequence models is the forget gate. While Transformers do not have an explicit recurrent form, we show that a forget gate can be naturally incorporated into Transformers by down-weighting the unnormalized attention scores in a data-dependent way. We name this attention mechanism Forgetting Attention and the resulting model the Forgetting Transformer (FoX). We show that FoX outperforms the Transformer on long-context language modeling, length extrapolation, and short-context downstream tasks, while performing on par with the Transformer on long-context downstream tasks. Moreover, it is compatible with the FlashAttention algorithm and does not require any positional embeddings. Several analyses, including the needle-in-the-haystack test, show that FoX also retains the Transformer’s superior long-context capabilities over recurrent sequence models such as Mamba-2, HGRN2, and DeltaNet. We also introduce a “Pro” block design that incorporates some common architectural components in recurrent sequence models and find it significantly improves the performance of both FoX and the Transformer. Our code is available at https://github.com/zhixuan-lin/forgetting-transformer.
@@ 1 Introduction
To address this question, we leverage an important fact: many recurrent sequence models with a forget gate can be written in a parallel linear attention form (Katharopoulos et al., 2020) analogous to softmax attention (Yang et al., 2023; Dao & Gu, 2024). In this parallel form, the forget gate mechanism translates into down-weighing the unnormalized attention scores in a data-dependent way. Our key insight is that this exact mechanism is also applicable to softmax attention. We name this attention mechanism Forgetting Attention and the resulting model the Forgetting Transformer (FoX).
@@ 3 Forgetting Transformer
Our proposed model, the Forgetting Transformer (FoX), features a modified softmax attention mechanism with a forget gate. We name this attention mechanism Forgetting Attention. Similar to the gated variant of linear attention introduced in the previous section, we first compute a scalar forget gate $f_{t}=\sigma({\bm{w}}_{f}^{\top}{\bm{x}}_{t}+b_{f})\in{\mathbb{R}}$ for each timestep $t$. The output of the attention is then
@@
where $F_{ij}=\prod_{l=j+1}^{i}f_{l}$ and $D_{ij}=\log F_{ij}=\sum_{l=j+1}^{i}\log f_{l}$. This can be written in matrix form:
@@
Besides its natural connection to gated linear attention, Forgetting Attention can also be seen as a data-dependent and learnable version of ALiBi (Press et al., 2021). ALiBi applies a data-independent bias $b_{ij}=-(i-j)m_{h}$ to the attention logits, where $m_{h}$ is a fixed slope specific to each head $h$. It is easy to show that ALiBi is equivalent to using a fixed, head-specific, and data-independent forget gate $f_{t}^{(h)}=\exp(-m_{h})$. In Section 4.5, we verify the superiority of data-dependent forget gates over ALiBi.
@@ Results
Figure 3: Visualization of the forget gate weight matrix ${\bm{F}}$ and the attention score matrix ${\bm{A}}$ from two heads in different layers. Since ${\bm{A}}$ is very sparse, we only show entries with scores larger than $0.1$. These results use FoX (Pro). More examples can be found in Appendix F.10.
@@ 6 Conclusion
We discuss several limitations of our work and potential future work. First, due to our limited computing resources, our main experiments only use models up to 760M parameters, 48B tokens, and a training context length of $16384$ tokens. Thus, an important direction for future work is to test FoX at larger scales. Second, we only consider causal sequence modeling in this work. It would be interesting to extend Forgetting Attention to the non-causal case. Finally, we could potentially prune computation (e.g., KV-cache eviction) adaptively based on the forget gate values, which may greatly reduce training and inference costs.
@@ Appendix A Detailed FoX (Pro) layer computation
where $D_{ij}=\sum_{l=j+1}^{i}\log f_{l}$ with $D_{ii}=0$ for any $i$. We then apply the output normalization, output gate, and the final output projection:
@@
For the multi-head case, each head $h$ maintains an independent copy of the parameters and computes its output sequence $({\bm{y}}_{i}^{(h)})_{i=1}^{L}$ independently. All RMSNorms are also applied independently to each head. [^4] For the QK-norm implementation, we accidentally shared a single set of $d_{\text{head}}$ RMSNorm scaling parameters across different heads in each layer for our experiments (normally there should be one set of $d_{\text{head}}$ parameters for each head). We have verified that this has no observable impact on performance. ${\bm{y}}_{i}^{(h)}$’s from different heads are then summed together to obtain the final output ${\bm{y}}_{i}$. Note that similar to the standard Transformer, even though the computation and parameters of different heads are conceptually independent, most computations can be implemented equivalently by properly splitting/concatenating the intermediate vectors/weight matrices of different heads.
@@ B.2 Model parameters, estimated FLOPs, and throughput
We report the number of (non-embedding) parameters, estimated FLOPs, and training throughput in Table 6. For the recurrent models, we estimate FLOPs using their recurrent form for simplicity. Note that the FlashAttention kernels for FoX (Pro), Transformer (Pro), and FoX (LLaMA) are implemented in Triton by us on top of Flag Attention (FlagOpen, 2023) without significant optimization, while Transformer (LLaMA) uses the official FlashAttention implementation (Dao, 2023)) in CUDA. Also, we did not focus on optimizing the efficiency of components such as QK-norm and KV-shift. We expect these four models to have similar throughput if they are properly optimized. Since we use a long training context length compared to model size, recurrent models have a significant advantage in theoretical FLOPs due to their linear complexity. Though an exact FLOPs-matched comparison would be interesting, it will require recalibrating the scaling law for the long-context setting and is beyond the scope of this work.
@@ Appendix D Data-independent forget gate initialization
To understand our initialization for the fixed forget gate and the data-independent forget gate, we first define a function $T(b)=\frac{1}{-\log\sigma(b)}$. This function is defined such that $\sigma(b)^{T(b)}=1/e$ is always true (i.e., $T(b)$ is the timesteps needed to achieve a decay of $1/e$). We then initialize $b^{(h)}=b^{(h)}_{\text{init}}$ such that $T(b^{(h)}_{(\text{init})})=\exp(\log T_{\min}+(\log T_{\max}-\log T_{\min})\frac{h-1}{H-1})$, where $T_{\min}$ and $T_{\max}$ are hyperparameters and $H$ is the number of heads. For example, if $h=4$ and $(T_{\min},T_{\max})=(2,128)$, then we have $(T(b^{(1)}_{\text{init}}),T(b^{(2)}_{\text{init}}),T(b^{(3)}_{\text{init}}),T(b^{(4)}_{\text{init}}))=(2,8,32,128)$. As mentioned in the main text, A fixed forget gate with $(T_{\min},T_{\max})$ is equivalent to ALiBi with a minimum slope $\frac{1}{T_{\max}}$ and a maximum slope $\frac{1}{T_{\min}}$. We also tested this initialization for the data-dependent forget gate but did not find it useful, so we simply initialize $\{b^{(h)}\}_{h=1}^{H}$ to zero for the data-dependent forget gate. For the data-independent and fixed forget gate, zero initialization performs extremely poorly.
@@ F.10 Additional visualization of forget gate and attention score matrices
Figure 28: Visualization of the attention score matrix ${\bm{A}}$ from 16 heads in 4 different layers. These results use FoX (Pro). Since ${\bm{A}}$ is very sparse, we only show entries larger than $0.1$.